OpenStudy (anonymous):

Taylor is going on a trip to Europe. She is bringing $500 US dollars. On the way over, she can exchange her US Dollars for Euros. The exchange rate is $1.30 USD = 1 EUR. How much money will she have to spend in Europe, according to their monetary system? A. $130 EURO B. $650 EURO C. $50 EURO D. $384.62 EURO

OpenStudy (anonymous):

monetary system ??

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

that's what the question says .

OpenStudy (anonymous):

i am in class xth don't know monetary system

OpenStudy (anonymous):

basically just like their money system and how their money works there .

OpenStudy (anonymous):

1 Eur= 1.3 US, 1US = 1/1.3 EURO SO 500 us = ? = (500*1)/1.3 euros

OpenStudy (anonymous):

as per their monetary system

OpenStudy (anonymous):

its 384

OpenStudy (anonymous):

i'm confused /:

OpenStudy (anonymous):

its simple just divide 500 dollars by 1.30

OpenStudy (anonymous):

384 EUROS AND 60 EURO CENTS

OpenStudy (anonymous):

OH ! thanks both of you !! i just divided it and i got ittt .

OpenStudy (anonymous):

@psi9epsilon its 384.62

OpenStudy (anonymous):

welcome